OpenStudy (anonymous):

Simplify the product using FOIL. (4x – 4)(3x – 4)? 12x² – 28x + 16 12x² – 4x – 16 12x² + 4x – 16 12x² + 28x + 16

OpenStudy (rulnick):

First one again.

OpenStudy (rulnick):

4x 3x = 12x^2 4x(-4) = -16x -4(3x) = -12x -4(-4) = 16 add em up to get 12 x^2 - 28x + 16

OpenStudy (anonymous):

(4x – 4)(3x – 4) 12x^2 - 28x + 16